OpenStudy (anonymous):

How do tan(17π / 12) in its reduced radical form?

OpenStudy (anonymous):

write \(\frac{17\pi}{12}=\frac{3\pi}{4}+\frac{2\pi}{3}\) and then use "addition angle formula"

OpenStudy (anonymous):

\[\left( \sqrt{3}+1 \right)\div \left( \sqrt{3-1} \right)\] I gave it a try...Is that right? in reduced radical form?